OpenStudy (warpedkitten):

Not sure how to solve. Julia uses the shadow of a flagpole to estimate its height. She stands so that the tip of her shadow coincides with the tip of the flagpole’s shadow, as shown. Julia is 5 feet tall. The distance from the flagpole to Julia is 28 feet, and the distance from the tip of the shadows to Julia is 7 feet. Calculate the height of the flagpole. a. 13 feet b. 980 feet c. 39 feet d. 25 feet

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

since the triangles are similar, we can set up this proportion \[\Large \frac{x}{28+7} = \frac{5}{7}\] solve for x to get your answer

OpenStudy (warpedkitten):

25?

jimthompson5910 (jim_thompson5910):

correct